OpenStudy (anonymous):

whats the cube of 2x+5

OpenStudy (ash2326):

Use this formula here \[(a+b)^3=a^3+b^3+3ab^2+3a^2b\] here a=2x b=5

OpenStudy (ash2326):

Can you try? @amandaschlor

OpenStudy (anonymous):

i tried i just dont get it.

OpenStudy (ghostfire):

Which part you don't get? for example \[(2x)^{3} = 2 \times2\times2\times x ^{3}\]

OpenStudy (ghostfire):

try to write out the entire equation first by plug 2x and 5 in the formula that @ash326 provided.

OpenStudy (anonymous):

but that equation is not one of the choices for the answer, its muitple choice and that is not an answer for that.

OpenStudy (ghostfire):

what did you get for the answer? and what are the multiple choice answers?

OpenStudy (anonymous):

the multiple choice answer are : 8x^3 + 60x^2 + 150x + 125 or 125x^3 + 150x^2 + 60x + 8 or 8x^3 + 60x^2 + 125x + 150 or 150x^3 + 125x^2 + 60x + 8 which one is it ?

OpenStudy (ghostfire):

what did you get?